Session: Nov 2000 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (G.Maccaferri at Sat Nov 18 11:43:16 MET 2000) Success

\(Other\) (A. Polatidis at Mon Nov 20 12:18:02 MET 2000) Success
Observations went OK. Readback tests showed lower than normal recording quality. we are investigating this problem.

Torun (K.Borkowski at Mon Nov 20 14:17:14 MET 2000) Success
Possible formatter clock offset by +1s.

Jodrell_Bank (Alastair Gunn at Mon Nov 20 15:17:32 MET 2000) Success
No known problems.

Effelsberg (A. Kraus at Mon Nov 20 16:06:09 MET 2000) Minor failures
Scans between UT 0430-0459 lost. Trouble with the telescope control computer.

Noto (C.Trigilio at Mon Nov 20 16:52:10 MET 2000) Minor failures
Lost 10 min from 06:29 to 06:40 for antenna problems.

Westerbork (Tony Foley at Wed Nov 22 09:19:32 MET 2000) Minor failures
Some EMI about 07:33.